# Fault tolerant Operations in Majorana-based Quantum Codes: Gates, Measurements and High Rate Constructions

### v1

@@ Abstract
Majorana-based quantum computation in nanowires and neutral atoms has gained prominence as a promising platform to encode qubits and protect them against noise. In order to run computations reliably on such devices, a fully fault-tolerant scheme is needed for state preparation, gates, and measurements. However, current fault-tolerant schemes have either been limited to specific code families or have not been developed fully. In this work, we develop a general framework for fault-tolerant computation with logical degrees encoded into Majorana hardware. We emphasize the division between even and odd Majorana codes and how it manifests when constructing fault tolerant gadgets for these families. We provide transversal constructions and supplement them with measurements to obtain several examples of fault tolerant Clifford gadgets. For the case of odd codes, we give a novel construction for gadgets using quantum reference frames, that allows to implement operations that are forbidden due to parity superselection. We also provide a fault-tolerant measurement scheme for Majorana codes inspired by Steane error correction, enabling state preparation, measurement of logical operations and error correction. We also point out a construction for odd Majorana codes with transversal T gates. Finally, we construct an asympotically good quantum LDPC Majorana code with qubit degrees of freedom. Our work shows that all necessary elements of fault-tolerant quantum computation can be consistently implemented in fermionic hardware such as Majorana nanowires and fermionic neutral atoms.
@@ I Introduction
Quantum computation is inherently susceptible to errors, and significant effort has gone into developing fault-tolerant schemes, all of which fundamentally rely on quantum error correction. However, a persistent challenge is that the physical error rates of many existing platforms remain above the fault-tolerance threshold. While ongoing advances in hardware — including improvements in coherence times and gate fidelities — aim to address this, an alternative and complementary approach is topological quantum computing, which seeks to build hardware with intrinsic protection against local errors. Majorana zero modes (MZM) are believed to form at superconductor-semiconductor interface, and while progress in the discovery of MZM’s has been challenging, there have been some promising efforts [1, 2] and many clear proposals on their implementation [3, 4]. These modes are robust to local decoherence, hence one can encode logical information into these modes [5]. Majorana zero modes also correspond to Ising anyons [6], and one can realize logical gates on the encoded data by topologically protected physical operations. However, Ising anyons are not universal and in order to do universal computing, other topologically unprotected schemes are required [7]. Moreover, existing schemes still suffer from other noise processes [8] and require active error correction [4].
@@
In the context of qubits based on Majorana nanowires, previous work [4] has focused on encoding qubits into small superconducting islands composed of two or three nanowires, and then concatenating these small Majorana codes with known qubit codes such as surface codes or Floquet codes. While these designs are valuable since they can use existing techniques of fault-tolerant operations on qubit codes, these leave the benefits general Majorana codes afford on the table as well. Moreover, generally qubit-based codes are only tailored to Pauli noise models and would not be able to protect against fermionic noise such as quasiparticle poisoning. In nanowire based systems, quasiparticle poisoning can also cause gates to fail and measurements to become faulty, so the reliance on topological protection to suppress errors, especially during measurements becomes precarious. Hence, creating a general toolkit of fault-tolerant gadgets for these devices that does not rely on physical means of protection is crucial.
Previous studies have developed fault-tolerant gate implementations for Majorana color codes that have utilized code deformations and lattice surgery techniques [12, 22, 13] supplemented with magic state distillation [23]. However, code deformations are typically much slower since they operate in the adiabatic limit, hence exploring other means of performing fault tolerant gates which includes transversal implementations and measurements is also extremely valuable.
@@ II.1 Majorana fermions
Similar to the Pauli group, the order of this group is $2^{2n+1}$. In MZM based architectures, each physical Majorana mode can be individually manipulated using gates, or Majorana bilinears can be measured. Hence logical states, commonly qubit states can be encoded into these modes and controlled. While for concreteness, our work will be mostly formulated in the language of Majorana zero modes realized in nanowires, there are other platforms in which Majorana computation can be realized.
@@
Similarly, electrons in dots and donors in silicon realize complex fermions and can be braided using tunneling [29] and pairing induced by the proximity effect with a superconductor. Moreover, qubit-fermion gates with hyperfine qubits in these platforms yield further flexibility [30, 27].
@@ II.5 Implementation of the $\mathrm{BRAID}_{2}$ and $\mathrm{BRAID}_{4}$ gates
Density-density interactions are implemented by using Rydberg interactions [38, 25]. For a general $\mathrm{BRAID}_{4}$ gate, we can conjugate 31 with $\mathrm{BRAID}_{2}$ gates to get the desired form.
@@ V Transversal gate gadgets
V.1 Detail of $\mathrm{BRAID}_{2}$ and $\mathrm{BRAID}_{4}$ tranversality
@@ V.2 Mode permutation to achieve logical gates
$G=\text{\footnotesize$\mathrm{CZ}$}(1,2)\text{\footnotesize$\mathrm{CNOT}$}(2\rightarrow 1)$ (69)
@@ X Majorana quantum LDPC code
Thus we have found a $N,\sqrt{2N},\sqrt{N/2}]$ code, which implies it is high rate and high distance, in fact, it has comparable parameters to a hypergraph product construction [47].
@@
of length $2^{n-1}$, where the ones vector is of size $2^{n-2}$. Combining 140 and 138 gives all ones in the right submatrix. Combining this with 137 gives the total parity $P_{\mathrm{tot}}$ inside the stabilizer, and concludes the proof.
@@
This gives an $8,4,2]]_{f}$ code, with eight physical fermions. The above construction gives logical qubits, but one can also find logical fermions by choosing an anti-commuting basis, although such logicals may not be of a CSS type.
@@ Appendix A Parity superselection as a constraint on operators
Conjugating $\rho=\rho_{e}$, with $U$, it is clear that one gets $\rho_{o}$ corresponds iff $U_{eo}$ is non-zero. However, parity SSR only allows $U$ to have a direct sum structure, so we arrive at a contradiction. This means that we cannot implement parity switching operations unitarily.
@@ Appendix D Encoding circuit with reference modes
To prepare the general state in 189, one can use the circuit

### v2

@@ Abstract
Majorana-based quantum computation in nanowires and neutral atoms has gained prominence as a promising platform to encode qubits and protect them against noise. In order to run computations reliably on such devices, a fully fault-tolerant scheme is needed for state preparation, gates, and measurements. However, current fault-tolerant schemes have either been limited to specific code families or have not been developed fully. In this work, we develop a general framework for fault-tolerant computation with logical degrees encoded into Majorana hardware. We emphasize the division between even and odd Majorana codes and how it manifests when constructing fault tolerant gadgets for these families. We provide transversal constructions and supplement them with measurements to obtain several examples of fault tolerant Clifford gadgets. For the case of odd codes, we give a novel construction for gadgets using quantum reference frames, that allows to implement operations that are forbidden due to parity superselection. We also provide a fault-tolerant measurement scheme for Majorana codes inspired by Steane error correction, enabling state preparation, measurement of logical operations and error correction. We also point out a construction for odd Majorana codes with transversal T gates. Finally, we construct a high rate quantum LDPC Majorana code with logical qubits. Our work shows that all necessary elements of fault-tolerant quantum computation can be consistently implemented in fermionic hardware such as Majorana nanowires and fermionic neutral atoms.
@@ I Introduction
Quantum computation is inherently susceptible to errors, and significant effort has gone into developing fault-tolerant schemes, all of which fundamentally rely on quantum error correction. However, a persistent challenge is that the physical error rates of many existing platforms remain above the fault-tolerance threshold. While ongoing advances in hardware — including improvements in coherence times and gate fidelities — aim to address this, an alternative and complementary approach is topological quantum computing, which seeks to build hardware with intrinsic protection against local errors. Majorana zero modes (MZM) are believed to form at superconductor-semiconductor interface, and while progress in the discovery of MZM’s has been challenging, there have been some promising efforts [1, 2] and many clear proposals on their implementation [3, 4]. These modes are robust to local decoherence, hence one can encode logical information into these modes [5]. Majorana zero modes also correspond to Ising anyons [6], and one can realize logical gates on the encoded data by topologically protected physical operations. However, Ising anyons are not universal and in order to do universal computing, other topologically unprotected schemes are required [7]. Moreover, existing measurement based schemes still suffer from other noise processes [8] and require active error correction [4].
@@
In the context of qubits based on Majorana nanowires, previous work [4] has focused on encoding qubits into small superconducting islands composed of two or three nanowires, and then concatenating these small Majorana codes with known qubit codes such as surface codes or Floquet codes. While these designs are valuable since they can use existing techniques of fault-tolerant operations on qubit codes, these leave the benefits general Majorana codes afford on the table as well. Moreover, generally qubit-based codes are only tailored to Pauli noise models and would not be able to protect against native fermionic noise. In nanowire-based systems, quasiparticle poisoning can also cause gates to fail and measurements to become faulty, so the reliance on topological protection to suppress errors, especially during measurements becomes precarious. Hence, creating a general toolkit of fault-tolerant gadgets for these devices that does not rely on physical means of protection is crucial.
Previous studies have developed fault-tolerant gate implementations for Majorana color codes that have utilized code deformations and lattice surgery techniques [12, 22, 13] supplemented with magic state distillation [23]. However, lattice surgery techniques require $d$ rounds of syndrome extraction, hence exploring other means of performing fault tolerant gates which includes transversal implementations and measurements is also extremely valuable.
@@ II.1 Majorana fermions
Similar to the Pauli group, the order of this group is $2^{2n+1}$. In MZM based architectures, each physical Majorana mode can be individually manipulated using gates, or Majorana bilinears can be measured. Hence logical states, commonly qubit states can be encoded into these modes and controlled.
@@
It is worth pointing out the normal physical setting for these systems corresponds to a particle number conservation constraint. More precisely, this means that physical Hamiltonian $H_{\text{f}}$ commutes with the total fermion number $N_{\text{tot}}$, which is a symmetry, and hence $H_{\text{f}}$ can be written in a block diagonal form, and similarly, the Hilbert space can also be written in this block diagonal basis
$\mathcal{H}_{f}=\bigoplus_{i=0}^{N}\mathcal{H}_{i}$ (7)
Moreover, the native unitaries are also considered to be number conserving operations. However, this is not an essential constraint, and people have shown proposals where the number conserving constraint can be lifted, either using reference systems [28] or by engineering specific pairing gates [27]. Doing so changes the symmetry from $S=N_{\text{tot}}$ to $S=P_{\text{tot}}$, so $H_{\text{f}}$ is instead block diagonal in the total parity basis, and the Hilbert space $\mathcal{H}_{f}$ also becomes a $Z_{2}$ graded space. For our purpose, this means that we can use Majorana modes to represent physical degrees of freedom instead of fermions, and apply parity-preserving unitaries to our system.
@@ II.5 Implementation of the $\mathrm{BRAID}_{2}$ and $\mathrm{BRAID}_{4}$ gates
Density-density interactions are implemented by using Rydberg interactions [38, 25]. For a general $\mathrm{BRAID}_{4}$ gate, we can conjugate 32 with $\mathrm{BRAID}_{2}$ gates to get the desired form.
@@ V Transversal gate gadgets
V.1 Detail of $\mathrm{BRAID}_{2}$ and $\mathrm{BRAID}_{4}$ transversality
@@ V.2 Mode permutation to achieve logical gates
$G=\text{\footnotesize$\mathrm{CZ}$}(1,2)\text{\footnotesize$\mathrm{CNOT}$}(2,1)$ (70)
@@ X Majorana quantum LDPC code
Thus we have found a $N,\sqrt{2N},\sqrt{N/2}]$ code, which implies it is high rate and high distance, in fact, the distance is comparable to a hypergraph product (HGP) construction [47], but $k$ is $O(\sqrt{N}$ as opposed to $O(N)$ for HGP codes. Moreover, check weights also grow as $O(\log(N))$, while for HGP codes, the checks are normally asssumed to be constant weight.
@@
of length $2^{n-1}$, where the ones vector is of size $2^{n-2}$. Combining 141 and 139 gives all ones in the right submatrix. Combining this with 138 gives the total parity $P_{\mathrm{tot}}$ inside the stabilizer, and concludes the proof.
@@
This gives an $8,4,2]]_{f}$ code, with eight physical fermions. The above construction gives logical qubits, but one can also find logical fermions by choosing an anti-commuting basis, although such logicals may not be of a CSS type. We note that there has been some subsequent work on Majorana LDPC codes that encode logical fermions [48].
@@ Appendix A Parity superselection as a constraint on operators
Conjugating $\rho=\rho_{e}$, with $U$, it is clear that one gets $\rho_{o}$ corresponds iff $U_{eo}$ is non-zero. However, parity SSR only allows $U$ to have a block diagonal structure, so we arrive at a contradiction. This means that we cannot implement parity switching operations unitarily.
@@ Appendix D Encoding circuit with reference modes
To prepare the general state in 190, one can use the circuit